10
Draw the graph of f(x) = 52-2 No + 9 8 7 6 5 4 3 2 1 5 6 7 8 9
Answer #1

- For y = f(x) = 54-2 First we should draw table xroro121314151 y = 5²4 525i 525 | 125 / i plot the points (0,5), (1,5) (2,4)een enorme 10 .- 3.5) 15 on x-axis-tacken X-values and on y-axis tacken values of f(x)=5^(x-2).

een enorme 10 .- 3.5) 15 on x-axis-tacken X-values and on y-axis tacken values of f(x)=5^(x-2).